### v2

@@ 2 Problem Statement and Related Literature
Random Utility Interpretation of S-RAM. The S-RAM is equivalent to the Marginal Distribution Model (MDM) proposed by Natarajan et al. (2009), an important class of semi-parametric choice models. Let $\tilde{\epsilon}_{1},\ldots,\tilde{\epsilon}_{n}$ be random variables with marginals $F_{1},\ldots,F_{n}.$ Let $\Theta$ denote the set of joint distributions for $\tilde{\boldsymbol{\epsilon}}$ with the given marginals $F_{1},F_{2},\ldots,F_{n}$.
@@
3 Step 1: Checking S-RAM Representability to The Data
@@
Under Assumptions 1 and 2, there exists an S-RAM to represent $\mathcal{P}^{\mathcal{K}}$ if and only if there exists $(\lambda_{i}^{k}\in\mathbb{R}:i\in[m],k\in\mathcal{K})$, such that for any distinct $k,l\in\mathcal{K}$ and $j\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)}$,
$\begin{aligned} &\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}<\sum_{i\in[m]}\lambda_{i}^{l}a_{ij}\quad\text{if}\quad p_{j}^{k}<p_{j}^{l},\\ &\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}=\sum_{i\in[m]}\lambda_{i}^{l}a_{ij}\quad\text{if}\quad 0<p_{j}^{k}=p_{j}^{l}<1.\end{aligned}$ (3)
@@
As a result, checking whether the given data $\mathcal{P}^{\mathcal{K}}$ satisfies the S-RAM hypothesis can be accomplished by solving a linear program with $\mathcal{O}(mK)$ continuous variables and $\mathcal{O}(nK)$ constraints.
@@
For any collection $\mathcal{I}^{\mathcal{K}},$ let $\mathcal{P}_{\textsc{S-RAM}}(\mathcal{K})$ denote the collection of probabilities for $\mathcal{I}^{\mathcal{K}}$, which are representable by any S-RAM. Due to the characterization in Theorem 3.1, we have the following succinct description for S-RAM: for any collection $\mathcal{I}^{\mathcal{K}}$, $\mathcal{P}_{\textsc{S-RAM}}(\mathcal{K})=\text{Proj}_{\boldsymbol{x}}(\Pi_{\mathcal{K}}):=\left\{\boldsymbol{x}:(\boldsymbol{x},\boldsymbol{\lambda})\in\Pi_{\mathcal{K}}\right\},$ where $\Pi_{\mathcal{K}}$ is defined as
$\Pi_{\mathcal{K}}=\Big\{(\boldsymbol{x},\boldsymbol{\lambda}):\,$ $\sum_{j\in\mathcal{I}^{(k)}}a_{ij}x_{j}\leq b_{i},\ \forall k\in\mathcal{K},\forall i\in[m],0\leq x_{j}\leq 1,\ \forall k\in\mathcal{K},\forall j\in\mathcal{I}^{(k)},$
@@
$\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}<\sum_{i\in[m]}\lambda_{i}^{l}a_{ij}\quad\text{if}\quad x_{j}^{k}<x_{j}^{l},\forall k,l\in\mathcal{K},\forall j\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)},$ (4)
@@
$\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}=\sum_{i\in[m]}\lambda_{i}^{l}a_{ij}\quad\text{if}\quad 0<x_{j}^{k}=x_{j}^{l}<1,\forall k,l\in\mathcal{K},\forall j\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)}\Big\}.$
4 Step 2: Nonparametric Predictions with A New Polytope
@@
When the data $\mathcal{P}^{\mathcal{K}}$ is S-RAM-representable, we proceed to Step 2 with the workflow in Figure 1. As an application of the exact characterization derived in Theorem 3.1, we develop a nonparametric data-driven approach for making predictions for a new structured polytope $\text{conv}(\mathcal{X}^{\text{new}})$ where $\mathcal{X}^{\text{new}}:=\{\boldsymbol{x}\in\{0,1\}^{n}\mid\boldsymbol{A}\boldsymbol{x}=\boldsymbol{b},x_{j}=0,\forall j\notin\mathcal{I}^{\text{new}}\}$ and $\mathcal{I}^{\text{new}}\notin\mathcal{I}^{\mathcal{K}}.$ The key idea of the nonparametric approach is as follows. To predict the value of a function on $\text{conv}(\mathcal{X}^{\text{new}})$, we consider all S-RAMs that are consistent with the observed data, and use the worst-case (or best-case) value of this function across the consistent models as our estimate. Thus, robust optimization forms the basis of our approach for allowing data to select a suitable S-RAM based on the prediction task at hand.
4.1 A Robust Optimization Formulation for Prediction
@@
The collection of all S-RAM choice probability vectors $\boldsymbol{x}=(x_{j}:j\in\mathcal{I}^{\text{new}})$ for the new polytope $\text{conv}(\mathcal{X}^{\text{new}})$ being consistent with the observed data $\boldsymbol{p}^{\mathcal{K}}$ is given by $\mathcal{U}_{\text{new}}:=\left\{\boldsymbol{x}:(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x})\in\mathcal{P}_{\textsc{S-RAM}}(\mathcal{K}^{\prime})\right\},$ where $\mathcal{I}^{\mathcal{K}^{\prime}}=\mathcal{I}^{\mathcal{K}}\cup\{\mathcal{I}^{\text{new}}\}$. Let $f(\boldsymbol{x})$ denote any function of interest for the prediction problem e.g., the expected revenue of a new assortment, the flow of a chosen link in a new network, total welfare of a new assignment, or other performance measures derived from the choice probabilities. Define the worst case estimate of $f(\boldsymbol{x})$ with $\boldsymbol{x}\in\mathcal{U}_{\text{new}}$ as $\underline{f}_{\text{new}}$. Due to the exact characterization of S-RAM in (3), we obtain the following formulation for the prediction problem with $\text{conv}(\mathcal{X}^{\text{new}})$.
@@
s.t. $x_{j}\ \geq\ p_{j}^{k}\quad\text{if}\quad\sum_{i\in[m]}\lambda_{i}^{\text{new}}a_{ij}\geq\sum_{i\in[m]}\lambda_{i}^{k}a_{ij},\quad\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$ (5a)
$x_{j}\ \leq\ p_{j}^{k}\quad\text{if}\quad\sum_{i\in[m]}\lambda_{i}^{\text{new}}a_{ij}\leq\sum_{i\in[m]}\lambda_{i}^{k}a_{ij},\quad\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$ (5b)
$\sum_{j\in\mathcal{I}^{\text{new}}}a_{ij}x_{j}\leq b_{i},\qquad\qquad\qquad\qquad\quad\quad\quad\quad\forall i\in[m],$
@@
4.2 A Compact-sized Integer Formulation for Prediction
@@
$\text{s.t.}\$- $\delta_{j}^{\text{new},k}\leq\sum_{i\in[m]}\lambda_{i}^{\text{new}}a_{ij}-\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}\leq 1-(1+\epsilon)\delta_{j}^{\text{new},k},\,\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$ (6a)
- $\delta_{j}^{k,\text{new}}\leq\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}-\sum_{i\in[m]}\lambda_{i}^{\text{new}}a_{ij}\leq 1-(1+\epsilon)\delta_{j}^{k,\text{new}},\;\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$ (6b)
$\delta_{j}^{\text{new},k}-1\ \leq\ p_{j}^{k}-x_{j}\ \leq\ 1-\delta_{j}^{k,\text{new}},\qquad\qquad\qquad\quad\;\;\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$ (6c)
- $(\delta_{j}^{\text{new},k}+\delta_{j}^{k,\text{new}})\ \leq\ p_{j}^{k}-x_{j}\ \leq\ \delta_{j}^{\text{new},k}+\delta_{j}^{k,\text{new}},\ \qquad\quad\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$ (6d)
@@
5 Step $1^{\prime}$: Nonparametric Estimation to Find the Best Fit
@@
5.1 A Nonparametric Estimation Formulation for S-RAM
Given observed data $\boldsymbol{p}^{\mathcal{K}}$, suppose that a loss function $\boldsymbol{x}^{\mathcal{K}}\mapsto\text{loss}(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x}^{\mathcal{K}})$ measures the degree of inconsistency in approximating the data $\boldsymbol{p}^{\mathcal{K}}$ with an S-RAM-consistent probability assignment $\boldsymbol{x}^{\mathcal{K}}\in\mathcal{P}_{\textsc{S-RAM}}(\mathcal{K}).$ We assume the loss function is non-negative and convex, and satisfies the property that $\text{loss}(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x}^{\mathcal{K}})=0$ if and only if $\boldsymbol{x}^{\mathcal{K}}=\boldsymbol{p}^{\mathcal{K}}$. Suppose that $(n_{k}:k\in\mathcal{K})$ is a vector of non-negative weights over polytopes in the historical dataset. Then common choices include a norm-based loss such as $\sum_{k\in\mathcal{K}}n_{k}\|\boldsymbol{p}_{k}-\boldsymbol{x}_{k}\|$ or a Kullback-Liebler divergence based loss such as $\sum_{k\in\mathcal{K}}n_{k}\sum_{i\in\mathcal{I}^{(k)}}p_{i}^{k}\log(x_{i}^{k}/p_{i}^{k})$. We define the limit of the S-RAM , denoted $\mathcal{L}(\boldsymbol{p}^{\mathcal{K}})$, is the minimum possible loss achievable by approximating $\boldsymbol{p}^{\mathcal{K}}$ using a model consistent with S-RAM:
$\mathcal{L}(\boldsymbol{p}^{\mathcal{K}})=\inf\left\{\text{loss}(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x}^{\mathcal{K}})\,:\,\boldsymbol{x}_{\mathcal{K}}\in\mathcal{P}_{\textsc{S-RAM}}(\mathcal{K})\right\}.$ (7)
From the above definition, evaluating the limit $\mathcal{L}(\boldsymbol{p}^{\mathcal{K}})$ amounts to finding a probability assignment $\boldsymbol{x}^{\mathcal{K}}_{\ast}$ in the S-RAM class that best fits the observed data $\boldsymbol{p}^{\mathcal{K}}$. Specifically, suppose we take $\text{loss}(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x}^{\mathcal{K}})=-\sum_{k\in\mathcal{K}}n_{k}\sum_{j\in\mathcal{I}^{(k)}}p_{j}^{k}\log(x_{j}^{k}/p_{j}^{k})$ with the weight $n_{k}$ for $k\in\mathcal{K}$ to be equal to the number of observations of a polytope $\text{conv}(\mathcal{X}^{(k)})$ in the dataset. A solution to (7) can be viewed as obtained from maximum likelihood estimation without any parametric assumption on S-RAM. Due to the characterization of S-RAM in Theorem 3.1, we have the equivalent formulation for $\mathcal{L}(\boldsymbol{p}^{\mathcal{K}})$ below.

### v3

@@ 2.3.3. Random Utility Interpretation of s-RAM
The s-RAM is equivalent to the Marginal Distribution Model (MDM) proposed by (Natarajan et al., 2009), an important class of semi-parametric choice models. Let $\tilde{\epsilon}_{1},\ldots,\tilde{\epsilon}_{n}$ be random variables with marginals $F_{1},\ldots,F_{n}.$ Let $\Theta$ denote the set of joint distributions for $\tilde{\boldsymbol{\epsilon}}$ with the given marginals $F_{1},F_{2},\ldots,F_{n}$.
@@
3. Step 1: Checking s-RAM Representability to The Data
@@
Under Assumptions 1 and 2, there exists an s-RAM to represent $\mathcal{P}^{\mathcal{K}}$ if and only if there exists $(\lambda_{i}^{k}\in\mathbb{R}:i\in[m],k\in\mathcal{K})$, such that
(3) $\begin{aligned} &\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}<\sum_{i\in[m]}\lambda_{i}^{l}a_{ij}\quad\text{if}\quad p_{j}^{k}<p_{j}^{l},\\ &\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}=\sum_{i\in[m]}\lambda_{i}^{l}a_{ij}\quad\text{if}\quad 0<p_{j}^{k}=p_{j}^{l}<1.\end{aligned}$
@@
for any distinct $k,l\in\mathcal{K}$ and $j\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)}$. As a result, checking whether the given data $\mathcal{P}^{\mathcal{K}}$ satisfies the s-RAM hypothesis can be accomplished by solving a linear program with $\mathcal{O}(mK)$ continuous variables and $\mathcal{O}(nK)$ constraints.
@@
For any collection $\mathcal{I}^{\mathcal{K}},$ let $\mathcal{P}_{\textsc{s-RAM}}(\mathcal{K})$ denote the collection of probabilities for $\mathcal{I}^{\mathcal{K}}$, which are representable by any s-RAM. Due to the characterization in Theorem 3.1, we have the following succinct description for s-RAM: for any collection $\mathcal{I}^{\mathcal{K}}$,
$\mathcal{P}_{\textsc{s-RAM}}(\mathcal{K})=\text{Proj}_{\boldsymbol{x}}(\Pi_{\mathcal{K}}):=\left\{\boldsymbol{x}:(\boldsymbol{x},\boldsymbol{\lambda})\in\Pi_{\mathcal{K}}\right\},$
where $\Pi_{\mathcal{K}}$ is defined as
@@
(4) $\Pi_{\mathcal{K}}=\left\{(\boldsymbol{x},\boldsymbol{\lambda}):\begin{aligned} &\sum_{j\in\mathcal{I}^{(k)}}a_{ij}x_{j}\leq b_{i},\ \forall k\in\mathcal{K},\forall i\in[m],\ 0\leq x_{j}\leq 1,\ \forall k\in\mathcal{K},\forall j\in\mathcal{I}^{(k)},\\ &\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}<\sum_{i\in[m]}\lambda_{i}^{l}a_{ij}\quad\text{if}\quad x_{j}^{k}<x_{j}^{l},\ \forall k,l\in\mathcal{K},\forall j\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)},\\ &\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}=\sum_{i\in[m]}\lambda_{i}^{l}a_{ij}\quad\text{if}\quad 0<x_{j}^{k}=x_{j}^{l}<1,\ \forall k,l\in\mathcal{K},\forall j\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)}\end{aligned}\right\}.$
While Theorem 3.1 provides a general and abstract characterization of s-RAM representability, it may be difficult to interpret directly. We therefore present the following examples to show how this characterization simplifies in concrete combinatorial problems. Taking the longest path problem as an example, the abstract ordering conditions reduce to intuitive comparisons of edge probabilities to be consistent with the relative shadow prices across instances, leading to simple linear feasibility checks. This example provides intuition for the characterization, which in turn facilitates understanding of the prediction and estimation formulations introduced later. For the representability conditions of the other examples listed in Table 1, we refer the reader to Section B.2 in the appendix.
Example 3.2 (s-RAM Characterization for the Longest Path Problem).
Suppose that $\boldsymbol{G}=(s\cup t\cup\boldsymbol{V},\boldsymbol{E})$ is the complete Directed Acyclic Graph (DAG) that describes all possible nodes and edges of the network, and we have flow data for a collection $K$ DAGs, denoted as $\boldsymbol{G}^{1},\boldsymbol{G}^{2},\cdots,\boldsymbol{G}^{K},$ with $\boldsymbol{G}^{k}=(s\cup t\cup\boldsymbol{V}^{k},\boldsymbol{E}^{k})$, for every $k\in\mathcal{K}$. For a graph $\boldsymbol{G}^{k}$, let $p_{ij}^{k}\in[0,1]$ denote the flow of edge $(i,j)$ in $\boldsymbol{G}^{k}$. For $k\in\mathcal{K}$, the convex hull of the longest path problem is given as
(5) $\operatorname{conv}(\mathcal{X}^{(k)})=\left\{\boldsymbol{x}:\begin{aligned} &\sum_{j:(i,j)\in\boldsymbol{E}^{k}}x_{ij}-\sum_{j:(j,i)\in\boldsymbol{E}^{k}}x_{ji}=\begin{cases}1,&i=s,\\ -1,&i=t,\\ 0,&i\in\boldsymbol{V}^{k},\end{cases}\\ &x_{ij}\geq 0,\quad\forall(i,j)\in\boldsymbol{E}^{k}\end{aligned}\right\}.$
Given observed flow data $\boldsymbol{p}^{\mathcal{K}}=(p_{ij}^{k}:(i,j)\in\boldsymbol{E}^{k},k\in\mathcal{K})$, by Theorem 3.1, the s-RAM characterization for longest path problems is given below.
Under Assumption 1 and 2, an observed flow collection $\boldsymbol{p}^{\mathcal{K}}$ is representable by an s-RAM if and only if there exists $(\lambda_{i}^{k}\in\mathbb{R}:i\in s\cup t\cup\boldsymbol{V}^{k},k\in\mathcal{K})$, such that for any two distinct graph $k,l\in\mathcal{K}$ containing a common edge $(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},$
(6) $\begin{aligned} &\lambda_{j}^{k}-\lambda_{i}^{k}>\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad p_{ij}^{k}>p_{ij}^{l},\\ &\lambda_{j}^{k}-\lambda_{i}^{k}=\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1.\end{aligned}$
@@
(7) $\begin{aligned} \max_{\epsilon,\boldsymbol{\lambda}}\quad&\epsilon\\ \text{s.t.}\quad&\lambda_{j}^{k}-\lambda_{i}^{k}\geq\lambda_{j}^{l}-\lambda_{i}^{l}+\epsilon\quad\text{if}\quad p_{ij}^{k}>p_{ij}^{l},\quad\forall k,l\in\mathcal{K},\forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\lambda_{j}^{k}-\lambda_{i}^{k}=\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1,\quad\forall k,l\in\mathcal{K},\forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l}.\end{aligned}$
4. Step 2: Nonparametric Predictions with A New Polytope
@@
When the data $\mathcal{P}^{\mathcal{K}}$ is s-RAM-representable, we proceed to Step 2 with the workflow in Figure 1. As an application of the exact characterization derived in Theorem 3.1, we develop a nonparametric data-driven approach for making predictions for a new structured polytope $\text{conv}(\mathcal{X}^{\text{new}})$ where $\mathcal{X}^{\text{new}}:=\{\boldsymbol{x}\in\{0,1\}^{n}\mid\boldsymbol{A}\boldsymbol{x}=\boldsymbol{b},x_{j}=0,\forall j\notin\mathcal{I}^{\text{new}}\}$ and $\mathcal{I}^{\text{new}}\notin\mathcal{I}^{\mathcal{K}}.$
The key idea of the nonparametric approach is as follows. To predict the value of a function on $\text{conv}(\mathcal{X}^{\text{new}})$, we consider all s-RAMs that are consistent with the observed data, and use the worst-case (or best-case) value of this function across the consistent models as our estimate. Thus, robust optimization forms the basis of our approach for allowing data to select a suitable s-RAM based on the prediction task at hand.
4.1. A Robust Optimization Formulation for Prediction
The collection of all s-RAM choice probability vectors $\boldsymbol{x}=(x_{j}:j\in\mathcal{I}^{\text{new}})$ for the new polytope $\text{conv}(\mathcal{X}^{\text{new}})$ being consistent with the observed data $\boldsymbol{p}^{\mathcal{K}}$ is given by
$\mathcal{U}_{\text{new}}:=\left\{\boldsymbol{x}:(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x})\in\mathcal{P}_{\textsc{s-RAM}}(\mathcal{K}^{\prime})\right\},$
@@ 4.1. A Robust Optimization Formulation for Prediction
where $\mathcal{I}^{\mathcal{K}^{\prime}}=\mathcal{I}^{\mathcal{K}}\cup\{\mathcal{I}^{\text{new}}\}$. Let $f(\boldsymbol{x})$ denote any function of interest for the prediction problem e.g., the expected revenue of a new assortment, the flow of a chosen link in a new network, total welfare of a new assignment, or other performance measures derived from the choice probabilities. Define the worst case estimate of $f(\boldsymbol{x})$ with $\boldsymbol{x}\in\mathcal{U}_{\text{new}}$ as $\underline{f}_{\text{new}}$. Due to the exact characterization of s-RAM in (4), we obtain the following formulation for the prediction problem with $\text{conv}(\mathcal{X}^{\text{new}})$.
@@
(8a) s.t. $x_{j}\ \geq\ p_{j}^{k}\quad\text{if}\quad\sum_{i\in[m]}\lambda_{i}^{\text{new}}a_{ij}\geq\sum_{i\in[m]}\lambda_{i}^{k}a_{ij},\quad\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$
(8b) $x_{j}\ \leq\ p_{j}^{k}\quad\text{if}\quad\sum_{i\in[m]}\lambda_{i}^{\text{new}}a_{ij}\leq\sum_{i\in[m]}\lambda_{i}^{k}a_{ij},\quad\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$
$\sum_{j\in\mathcal{I}^{\text{new}}}a_{ij}x_{j}\leq b_{i},\qquad\qquad\qquad\qquad\quad\quad\quad\;\,\forall i\in[m],$
@@
4.2. A Compact-sized Integer Formulation for Prediction
@@
(9a) $\text{s.t.}\$- $\delta_{j}^{\text{new},k}\leq\sum_{i\in[m]}\lambda_{i}^{\text{new}}a_{ij}-\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}\leq 1-(1+\epsilon)\delta_{j}^{\text{new},k},\,\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$
(9b) - $\delta_{j}^{k,\text{new}}\leq\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}-\sum_{i\in[m]}\lambda_{i}^{\text{new}}a_{ij}\leq 1-(1+\epsilon)\delta_{j}^{k,\text{new}},\;\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$
(9c) $\delta_{j}^{\text{new},k}-1\ \leq\ p_{j}^{k}-x_{j}\ \leq\ 1-\delta_{j}^{k,\text{new}},\qquad\qquad\qquad\quad\;\;\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$
(9d) - $(\delta_{j}^{\text{new},k}+\delta_{j}^{k,\text{new}})\ \leq\ p_{j}^{k}-x_{j}\ \leq\ \delta_{j}^{\text{new},k}+\delta_{j}^{k,\text{new}},\ \qquad\quad\forall k\in\mathcal{K},\forall j\in\mathcal{I}^{\text{new}}\cap\mathcal{I}^{(k)},$
@@
5. Step $1^{\prime}$: Nonparametric Estimation to Find the Best Fit
@@
5.1. A Nonparametric Estimation Formulation for s-RAM
Given observed data $\boldsymbol{p}^{\mathcal{K}}$, suppose that a loss function $\boldsymbol{x}^{\mathcal{K}}\mapsto\text{loss}(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x}^{\mathcal{K}})$ measures the degree of inconsistency in approximating the data $\boldsymbol{p}^{\mathcal{K}}$ with an s-RAM-consistent probability assignment $\boldsymbol{x}^{\mathcal{K}}\in\mathcal{P}_{\textsc{s-RAM}}(\mathcal{K}).$ We assume the loss function is non-negative and convex, and satisfies the property that $\text{loss}(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x}^{\mathcal{K}})=0$ if and only if $\boldsymbol{x}^{\mathcal{K}}=\boldsymbol{p}^{\mathcal{K}}$. Suppose that $(n_{k}:k\in\mathcal{K})$ is a vector of non-negative weights over polytopes in the historical dataset. Then common choices include a norm-based loss such as $\sum_{k\in\mathcal{K}}n_{k}\|\boldsymbol{p}_{k}-\boldsymbol{x}_{k}\|$ or a Kullback-Liebler divergence based loss such as $\sum_{k\in\mathcal{K}}n_{k}\sum_{i\in\mathcal{I}^{(k)}}p_{i}^{k}\log(x_{i}^{k}/p_{i}^{k})$.
We define the limit of the s-RAM , denoted $\mathcal{L}(\boldsymbol{p}^{\mathcal{K}})$, is the minimum possible loss achievable by approximating $\boldsymbol{p}^{\mathcal{K}}$ using a model consistent with s-RAM:
(10) $\mathcal{L}(\boldsymbol{p}^{\mathcal{K}})=\inf\left\{\text{loss}(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x}^{\mathcal{K}})\,:\,\boldsymbol{x}_{\mathcal{K}}\in\mathcal{P}_{\textsc{s-RAM}}(\mathcal{K})\right\}.$
From the above definition, evaluating the limit $\mathcal{L}(\boldsymbol{p}^{\mathcal{K}})$ amounts to finding a probability assignment $\boldsymbol{x}^{\mathcal{K}}_{\ast}$ in the s-RAM class that best fits the observed data $\boldsymbol{p}^{\mathcal{K}}$. Specifically, suppose we take $\text{loss}(\boldsymbol{p}^{\mathcal{K}},\boldsymbol{x}^{\mathcal{K}})=-\sum_{k\in\mathcal{K}}n_{k}\sum_{j\in\mathcal{I}^{(k)}}p_{j}^{k}\log(x_{j}^{k}/p_{j}^{k})$ with the weight $n_{k}$ for $k\in\mathcal{K}$ to be equal to the number of observations of a polytope $\text{conv}(\mathcal{X}^{(k)})$ in the dataset. A solution to (10) can be viewed as obtained from maximum likelihood estimation without any parametric assumption on s-RAM.
Due to the characterization of s-RAM in Theorem 3.1, we have the equivalent formulation for $\mathcal{L}(\boldsymbol{p}^{\mathcal{K}})$ below.